1. As you prepare to enter the workforce, you would like to start a systematic savings...

1.

As you prepare to enter the workforce, you would like to start a systematic savings plan in order to have enough money to retire after 30 years of employment.

Assume that your employers throughout your career pay annual bonuses. You have decided that you will put $ 8,000 per year into your savings account. You will make a deposit on the same date each year for 30 years. Assuming an interest rate of 9 %, how much will you have in your savings account at retirement?

2.

You had good intentions to start your savings plan when you started your career. However, you decided you needed a new car and needed to furnish your new condo with all of the right "stuff" instead. Ten years went by before you started your savings plan. Assuming the same facts in question # 11 except you only have 20 years to save instead of 30 years. How much will you have in your savings account at retirement?
Answer #1

1, Amount to be deposited each year = $8,000

Years of deposits = 30

Interest Rate = 9%

Total savings account amount at retirement = 8,000*30 = $240,000

Interest amount = 240,000*9% = $21,600

Total amount of savings account at retirement = 240,000+21,600 = $ 261,600

2, Amount to be deposited each year = $8,000

Years of deposits = 20

Interest Rate = 9%

Total savings account amount at retirement = 8,000*20 = $160,000

Interest amount = 160,000*9% = $14,400

Total amount of savings account at retirement = 160,000+14,400 = $ 174,400